Solve by factoring and using the principle of zero products.
25y^2 + 30y + 8 = 0

Answer by htmentor(1343)

Solve by factoring and using the principle of zero products.
25y^2 + 30y + 8 = 0
This can be factored as (5y+2)(5y+4)
For the expression to equal 0, either 5y+2 = 0 or 5y+4 = 0
Solving for y gives the results y = -2/5, y = -4/5